I believe I have the right answers just need help showing it step by step.
12.    Work backwards to write a quadratic equation that will have solutions of x = -1/2 and x = 4.  (Your equation must only have integer coefficients, meaning no fractions or decimals.)
13.   Write a quadratic equation that will have a solution of only x = 0. Note: this means there will be a double solution of x = 0.
 
y = x^2
 
14.   Write a quadratic equation that cannot be factored.
 
y = x^2 + 1
 
15.   The product of two consecutive positive integers is 72. Find the integers.
 
Two consecutive positive integers are 8 and 9 both have a product of 72.

Found 2 solutions by dkppathak, Edwin McCravy:
Answer by dkppathak(439)

Work backwards to write a quadratic equation that will have solutions of x = -1/2 and x = 4. (Your equation must only have integer coefficients, meaning no fractions or decimals.)
13. Write a quadratic equation that will have a solution of only x = 0. Note: this means there will be a double solution of x = 0.

y = x^2

14. Write a quadratic equation that cannot be factored.

y = x^2 + 1

15. The product of two consecutive positive integers is 72. Find the integers.

Two consecutive positive integers are 8 and 9 both have a product of 72.
Solution
12
x=1/2 and x+4
factors will be 2x-1 and another x-4
equation will be (2x-1)(x-4)=2x^2-9x+4=0
13. first factors and second factors are x
quadratic equation x*x=x^2
14. X^2+1 is correct this can not be factories for real factors but can be factories for imaginary or non real factors
15. let consecutive integers are X+1 and X
(x)(X+!)=72
X^2+x-72=0 the factors will be (x-9)(X-8)=0
integers will be 9 and 8

Answer by Edwin McCravy(20054)

I believe I have the right answers just need help showing it step by step.
12. Work backwards to write a quadratic equation that will have solutions of x = -1/2 and x = 4. (Your equation must only have integer coefficients, meaning no fractions or decimals.)

13. Write a quadratic equation that will have a solution of only x = 0.
Note: this means there will be a double solution of x = 0.
Do it the same way:


 
14. Write a quadratic equation that cannot be factored.
x²+x+1=0
 
15. The product of two consecutive positive integers is 72. Find the integers.
Let the consecutive integers be n and n+1
 
  

Since the problem says "POSITIVE" we can ignore the negative answer.
n = 8 is the smaller integer, so the larger is 8+1 or 9.
Answer: 8,9 

[You may wonder why I didn't just factor instead of using the quadratic
formula.  The reason was that in order to factor, you have to already 
be able to solve the problem in your head in order to do the factorization. 
So I reasoned that if you already have to know the answers in order to be 
able to find the answers, which are 8 and 9, then you're wasting your time
solving it. So I did it the only way that doesn't require that you must
have already solved the problem in order to solve the problem.] 
J

Edwin